Youve made the finals of the Science Olympics! As one of your tasks, youre given 7.0 g of copper and asked to make a wire,
Answer #1

given

mass m = 7*10^-3 kg

resistance R = 9 ohm

density d = 8900 kg/m^3

resistivity = \rho = 1.7*10^-8 ohm- m

A)

from R = \rho l/A

A = \rho l/R

density d = m/lA

d = m/l*\rhol/R = mR/\rhol^2

l = sqrt(7*10^-3*9/1.7*10^-8*8900) = 20.41 m

B)

area A = 1.7*10^-8*20.41/9 = 0.486*10^-8 m^2

A = 3.14*r^2 = 0.486*10^-8

radius r = 0.393*10^-4 m

diameter = 2*r = 0.786*10^-4 m

l = length =